# Understanding Performance of Long-Document Ranking Models through Comprehensive Evaluation and Leaderboarding

### v3

@@ B.2.2 Model Traning
  • In each column we show a relative gain over model’s respective FirstP baseline: The last column shows the average relative gain over FirstP. Best numbers are in bold: Our results are averaged over three seeds (but not necessarily prior art).
@@
All MS MARCO models were trained from scratch. Then these models were fine-tuned on Robust04. Note that except for the aggregation Transformers and TinyLLAMA, we use a base, i.e., a 12-layer Transformer Vaswani et al. (2017) models. TinyLLAMA has 22 layers and about 1B parameters. BERT-base is more practical then a 24-layer BERT-large and performs at par with BERT-large on MS MARCO and Robust04 Hofstätter et al. (2020a); Li et al. (2024). In our own experiments, we see that large (24 and more layers) model perform much better on the MS MARCO Passage collection, but we were not able to outperform 12-layer models on the MS MARCO Documents collection. Note that Longformer Beltagy et al. (2020), BigBird Zaheer et al. (2020), and DEBERTA base He et al. (2021), JINA Günther et al. (2023), and MOSAIC Portes et al. (2023) all have 12 layers, but a larger embedding matrix.
@@
One training epoch consisted in iterating over all queries and sampling one positive and one negative example with a subsequent computation of a pairwise margin loss. We used the minibatch size one with gradient accumulation over 16 steps. The learning rates are provided in the model configuration files (in the on-line repository). [^9] https://anonymous.4open.science/r/long_doc_rank_model_analysis_v2-78E9/. We used the AdamW optimizer Loshchilov and Hutter (2017) and a constant learning rate with a 20% linear warm-up Mosbach et al. (2020).
@@
We have learned that—unlike neural retrievers—cross-encoding rankers Nogueira and Cho (2019) are relatively insensitive to learning rates, their schedules, and the choice of loss functions. We were sometimes able to achieve better results using multiple negatives per query and a listwise margin loss (or cross-entropy). However, the gains were small and not consistent compared to a simple pairwise margin loss used in our work (in fact, using a listwise loss function sometimes lead to overfitting). Note again that this is different from neural retrievers where training is difficult without using a listwise loss and/or batch-negatives Karpukhin et al. (2020); Xiong et al. (2021); Qu et al. (2021); Zerveas et al. (2021); Formal et al. (2021).
@@
For MS MARCO, all models except PARADE-Transf-Pretr-LATEIR-L6 and PARADE-Transf-RAND-L2 were trained for one epoch: Further training did not improve (and sometimes degraded) accuracy. However, PARADE-Transf-RAND-L2 and PARADE-Transf-Pretr-LATEIR-L6 required two-to-three epochs to reach the maximum accuracy. In the case of Robust04, each model was finetuned for 100 epochs, but all epochs were short, so the overall training and evaluation time was comparable to that of fine-tuning on MS MARCO for a single epoch.
To reproduce our main results, we estimate that one needs about 6400 GPU hours: 6000 hours using NVIDIA A10 (or RTX 3090) with 24 GB of memory and 400 hours using NVIDIA A6000 with 48 GB of memory. A6000 was required only for TinyLLAMA.
@@
From our experience models trained on MS MARCO v2 performed worse on TREC 2021 queries compared to models trained on MS MARCO v1. This may indicate that models somehow learn to distinguish between original MS MARCO v1 documents and newly added ones (which did not have positive judgements in the training sets). As a result, these models are biased and tend to not rank these new documents well even when they are considered to be relevant by NIST assessors. For this reason, we used MS MARCO v2 data in a zero-shot transfer mode where ranking models trained on MS MARCO v1 are evaluated on TREC DL 2021 queries.
Table 8: Effectiveness of the PARADE Attention Model for Different Input Truncation Thresholds
@@
PARADE Attn (2 chunks) 0.408a $(+1.8\%)$ 0.653a $(+2.7\%)$ 0.499a $(+4.9\%)$ 0.544a $(+3.3\%)$ +$3.2\%$
PARADE Attn (3 chunks) 0.406a $(+1.3\%)$ 0.648a $(+1.7\%)$ 0.505a $(+6.1\%)$ 0.557a $(+5.7\%)$ +$3.7\%$
@@
PARADE Attn (5 chunks) 0.409a $(+2.0\%)$ 0.652a $(+2.4\%)$ 0.502a $(+5.6\%)$ 0.556a $(+5.5\%)$ +$3.9\%$
PARADE Attn (6 chunks) 0.411a $(+2.4\%)$ 0.653a $(+2.6\%)$ 0.504a $(+5.9\%)$ 0.554a $(+5.2\%)$ +$4.0\%$
B.2.3 Miscellaneous Notes
@@
To enable efficient training and evaluation of the large number of models, documents were truncated to have at most 1431 BERT tokens. In § B.3 (see Table 8) we show that for our top-performing model PARADE Attention Li et al. (2024) using a larger number of chunks only marginally improves outcomes. Depending on a dataset, the highest accuracy is achieved using either three or four chunks.
For SplitP approaches, queries were padded to 32 BERT tokens with padding being masked out during training (longer queries were truncated). For SplitP models with greedy partitioning into disjoint chunks, long document were split into at most three chunks containing 477 document tokens (each concatenated with up to 32 query tokens plus three special tokens).
@@
We evaluated 20+ models, but we had to exclude two LongT5 variants Guo et al. (2022) and RoFormer (with ROPE embeddings) Su et al. (2024) due to poor convergence and/or crashes. Specifically, even after 10 epochs of training LongT5 models were $\approx 10\%$ less accurate than BERT-base FirstP trained for one epoch. Given the uncertainty regarding the possible convergence of models as well as the need to train these for three epochs, we have to abandon this experiment as overly expensive. RoFormer models were failing due to CUDA errors when the context length exceeded 512: We were not able to resolve this issue.
B.3 Varying the Number of Chunks
@@
To understand if truncating input to have at most 1431 BERT tokens negatively affected model performance, we carried out an ablation study where one of the top-performing models was trained and evaluated using inputs of varying maximum lengths. To this end we used PARADE Attention with the number of input chunks varying from one to six. In that the same number of chunks was used during training and evaluation, i.e., we had to carry out six experiments. Similar to our main experiments, we trained each model using three seeds. We carried out this ablation experiment using our MS MARCO and Robust04 datasets.
@@
The results are presented in Table 8: We can see that—depending on the dataset—three or four input chunks are optimal. However, the additional gains over the FirstP baseline are at most 0.6% when averaged over all test sets.
@@
Gao and Callan 2022 carried out a similar ablation using ClueWeb09: Increasing the number of input chunks from three to six lead to only about 2.3% relative improvement in NDCG@20. However, even this modest gain could have been slightly inflated due to model not being trained directly on shorter inputs. Indeed, truncation of an input for a six-chunk model to one chunk is potentially less effective than training and evaluating the model using one-chunk data.
B.4 Reproducibility and Backbone Selection for SplitP Models
Table 9: Ranking Performance on MS MARCO and TREC DL.
@@ B.4 Reproducibility and Backbone Selection for SplitP Models
FirstP (E5-4K) zero-shot 0.380 0.641 0.722 0.317
@@
LongP (E5-4K) zero-shot 0.353a $(-7.1\%)$ 0.649 $(+1.3\%)$ 0.724 $(+0.3\%)$ 0.323 $(+1.8\%)$
@@
Table 10: Ranking Performance on Robust04.

### v4

@@ A.1.2 Model Training
  • In each column we show a relative gain over model’s respective FirstP baseline: The last column shows the average relative gain over FirstP. Best numbers are in bold: Our results are averaged over three seeds. Prior-art seed strategy is unknown.
@@
Table 5: Comparison between long-document models and respective FirstP (truncation) baselines for several backbone Transformer models as well as comparison to prior art. Results for MS MARCO, TREC DL, and Robust04.
All MS MARCO models were trained from scratch. Afterward, these models were fine-tuned on Robust04. Note that except for the aggregator Transformers (see D.2 for architectural details) and TinyLLAMA, we use a base, i.e., a 12-layer Transformer Vaswani et al. (2017) models. TinyLLAMA has 22 layers and about 1B parameters.
@@
BERT-base is more practical than a 24-layer BERT-large and performed on par with BERT-large on MS MARCO Documents and Robust04 Hofstätter et al. (2020a); Li et al. (2024). In our own preliminary experiments, we observed that the 24-layer BERT-large model performed much better on the MS MARCO Passage collection, but we were not able to outperform 12-layer BERT-base models on the MS MARCO Documents collection. Note that Longformer Beltagy et al. (2020), BigBird Zaheer et al. (2020), DEBERTA-base He et al. (2021), JINA Günther et al. (2023), and MOSAIC Portes et al. (2023) all have 12 layers, but a larger embedding matrix compared to BERT-base.
@@
One training epoch consisted in iterating over all queries and sampling one positive and one negative example with a subsequent computation of a pairwise margin loss. We used the mini-batch size one with gradient accumulation over 16 steps. The learning rates are provided in the model configuration files (in the code repository). [^7] https://github.com/searchivarius/long_doc_rank_model_analysis_v2/. We used the AdamW optimizer Loshchilov and Hutter (2017) and a constant learning rate with a 20% linear warm-up Mosbach et al. (2020).
@@
We have learned that—unlike neural retrievers—cross-encoding rankers Nogueira and Cho (2019) are relatively insensitive to learning rates, their schedules, and the choice of loss functions. We were sometimes able to achieve better results using multiple negatives per query and a listwise margin loss (or cross-entropy). However, the gains were small and not consistent compared to a simple pairwise margin loss used in our work (in fact, using a listwise loss function sometimes led to overfitting). Note again that this is different from neural retrievers where training is sometimes difficult without using a listwise loss and/or batch-negatives Karpukhin et al. (2020); Xiong et al. (2021); Qu et al. (2021); Zerveas et al. (2021); Formal et al. (2021).
@@
For MS MARCO, all models except PARADE-Transf-Pretr-LATEIR-L6 and PARADE-Transf-RAND-L2 were trained for one epoch: Further training did not improve (and sometimes degraded) accuracy. However, PARADE-Transf-RAND-L2 and PARADE-Transf-Pretr-LATEIR-L6 required two-to-three epochs to reach the maximum accuracy. For training with the debiased MS MARCO, we used only one epoch. In the case of Robust04, each model was fine-tuned for 100 epochs, but all epochs were short, so the overall training and evaluation time was comparable to that of fine-tuning on MS MARCO for a single epoch.
Except 1B-parameter TinyLLAMA Zhang et al. (2024), each model has about 100M parameters (see Table 11 for details). Despite training and testing 20+ models with three seeds, we estimate to have spent only about 6400 GPU hours for our main experiments. 96% of the time we used NVIDIA A10 (or similarly-powerful) RTX 3090 GPUs and 4% of the time we used NVIDIA A6000.
We believe this is roughly equivalent to pre-training a single 1B-parameter TinyLLAMA model with a next-token prediction objective, which required about 3400 GPU hours using a more powerful NVIDIA A100. This, in turn, this is only a tiny fraction of compute required to train LLAMA2 models (2% compared to a 7B LLAMA2 model). [^8] https://github.com/microsoft/Llama-2-Onnx/blob/main/MODEL-CARD-META-LLAMA-2.md
@@
From our experience, models trained on MS MARCO v2 performed worse on TREC 2021 queries compared to models trained on MS MARCO v1. This may indicate that models implicitly learned to distinguish between original MS MARCO v1 documents and newly added ones (which did not have positive judgements in the training sets). As a result, these models are biased and tend to rank newly added documents poorly even when they are considered to be relevant by NIST assessors. For this reason, we used MS MARCO v2 data only in a zero-shot transfer mode. To this end, ranking models trained on MS MARCO v1 were evaluated using TREC DL 2021 queries.
@@
PARADE Attn (2 chunks) 0.408a $(+1.8\%)$ 0.653a $(+2.7\%)$ 0.499a $(+4.9\%)$ 0.544a $(+3.3\%)$ $3.2\%$
PARADE Attn (3 chunks) 0.406a $(+1.3\%)$ 0.648a $(+1.7\%)$ 0.505a $(+6.1\%)$ 0.557a $(+5.7\%)$ $3.7\%$
@@
PARADE Attn (5 chunks) 0.409a $(+2.0\%)$ 0.652a $(+2.4\%)$ 0.502a $(+5.6\%)$ 0.556a $(+5.5\%)$ $3.9\%$
PARADE Attn (6 chunks) 0.411a $(+2.4\%)$ 0.653a $(+2.6\%)$ 0.504a $(+5.9\%)$ 0.554a $(+5.2\%)$ $4.0\%$
Table 6: Effectiveness of the PARADE Attention model for different input truncation thresholds. Results for MS MARCO, TREC DL, and Robust04.
A.1.3 Miscellaneous Notes
To enable efficient training and evaluation of the large number of models, both Robust04 and original MS MARCO documents were truncated to have at most 1431 BERT tokens. Thus, for SplitP approaches, queries were padded to 32 BERT tokens and long documents were split into at most three chunks, each containing 477 document tokens. However, for training on debiased MS MARCO, the truncation threshold was much higher: 8109 tokens.
@@ A.1.3 Miscellaneous Notes
In § B.2.1 (see Table 6) we show that for our top-performing model PARADE Attention Li et al. (2024) using a larger number of chunks only marginally improves outcomes. Depending on a dataset, the highest accuracy is achieved using either three or four chunks.
@@
We evaluated over 20 models, but we had to exclude two LongT5 variants Guo et al. (2022) and RoFormer (with ROPE embeddings) Su et al. (2024) due to poor convergence and/or crashes. Specifically, even after 10 epochs of training LongT5 models were $\approx 10\%$ less accurate than BERT-base FirstP trained for one epoch. Given the uncertainty regarding the possible convergence of models as well as the need to train these for three epochs, we had to abandon this experiment as overly expensive. RoFormer models were failing due to CUDA errors when the context length exceeded 512: We were not able to resolve this issue.
For bias-mitigation experiments, for several reasons, we used only a subset of models. First, we had to exclude all LongP models since none of them supported a context longer than 8192 tokens. In contrast, we trained our chunk-and-aggregate models on documents limited to 8109 tokens and then extrapolated their use to rank documents up to 32768 tokens long.
Second, we chose representative models with vastly different generalization properties. MaxP and PARADE Attention models performed well on MS MARCO FarRelevant in the zero-shot setting, but did not benefit much from in-domain fine-tuning. PARADE Transformer MRR dropped from 0.433 to 0.229 in the zero-shot setting, but increased up to 0.432 after in-domain fine-tuning. CEDR-KNRM also benefited a lot from fine-tuning on MS MARCO FarRelevant, but its zero-shot performance was at the level of the random-baseline.
Appendix B Additional Experimental Results
This section presents supplementary experimental results. In particular, we compute additional effectiveness metrics for MS MARCO, TREC DL, and Robust04. MS MARCO and TREC DL results are shown in Table 9. Results for Robust04 datasets are presented in Table 10. Evaluation results for rankers trained on the debiased MS MARCO and tested on short-document collections are shown in Table 3. Furthermore, we provide detailed results for MS MARCO FarRelevant in Table 4.
To justify the use of BERT-base Devlin et al. (2019) as the main model backbone, we carry out an ablation study in § B.1. Moreover, in this section we also compare our results to prior art, with the objective to boost trustworthiness of our experiments.
This is followed by a miscellaneous ablations subsection (§ B.2), where we carry out an efficiency evaluation and assess if truncating input to at most 1431 BERT tokens affected our experimental outcomes.
B.1 Backbone Selection for SplitP Models and Prior Art Comparison
B.1.1 Choice of a Backbone
To understand if using BERT-base as a backbone model for various SplitP (i.e., chunk-and-aggregate) approaches diminished models’ ability to process and understand long contexts, we carried out a focused comparison of several backbone models, including ELECTRA Clark et al. (2020) and DEBERTA He et al. (2021).
To this end, we used two methods: PARADE Li et al. (2024) Attention and MaxP. PARADE Attention model achieved the largest average gain over FirstP in our main experiments (see Table 1), whereas MaxP models were extensively benchmarked in the past Li et al. (2024); Dai and Callan (2019); Zhang et al. (2021).
Although prior work found ELECTRA to be a better backbone model in terms of absolute accuracy Li et al. (2024); Zhang et al. (2021), we found no systematic evaluation of the relationship between a backbone model and gains achievable over FirstP.
Results in Tables 1 and 5 confirm the overall superiority of both ELECTRA and DEBERTA over BERT-base. In that, DEBERTA models are nearly always more effective compared to ELECTRA models with biggest differences on Robust04. However, their relative effectiveness with respect to their respective FirstP baselines does not exceed that of BERT-base.
The same is true for LongP models. Except Longformer, they performed equally or worse compared to FirstP on 8 test sets out of 18. Moreover, all LongP models achieved lower average gains over FirstP (see the last column in Table 1). We conclude that to measure capabilities of chunk-and-aggregate or long-context models to understand and incorporate long documents, it appears to be beneficial to use BERT-base instead of ELECTRA or DEBERTA.
Finally, we would like to note that on standard benchmarks Big-Bird’s Zaheer et al. (2020) FirstP version always outperformed its LongP version (sometimes by as much as 10-15%), which seems to be puzzling. We noticed, however, that for shorter inputs, the model turns off sparse attention and prints the respective warning. Thus, we hypothesize that it is the use of sparse attention that causes this degradation. In contrast, the sparse attention implementation of the Longformer Beltagy et al. (2020) does not exhibit such a degradation (although with Longformer, not all attention is sparse: query-to-document attention is full). Although Big-Bird underperforms on standard benchmarks, after fine-tuning, it still matches the Longformer accuracy on MS MARCO FarRelevant (see Table 4).
B.1.2 Comparison to Prior Art
Note that one should not expect identical results due to differences in training regimes and candidate generators. In particular, in the case of Robust04, Li et al. 2024 used RM3, which is BM25 with pseudo-relevance feedback Jaleel et al. (2004). On Robust04, it is more effective than BM25 Robertson (2004), which we used as the first-stage retriever.
Another important comparison point is Robust04 results by Zhang et al. 2021 who were able to reproduce original MaxP results by Dai and Callan (2019), who used BERT-base as a backbone. In addition, they experimented with ELECTRA models “pre-finetuned” on MS MARCO. When comparing results for the BERT-base backbone, Zhang et al. (2021) have the maximum relative gain of 6.6% compared to ours 3.3%. However, in absolute terms we got higher NDCG@20 for both FirstP and MaxP. Their MaxP (ELECTRA) has comparable performance to ours on TREC DL 2019-2020, but it is 2-4% better on Robust04. In turn, our MaxP (DEBERTA) is better by 2-6%. Although we do not always match prior art using the same backbone models, we generally match or outperform prior results, which, we believe, boosts the trustworthiness of our experiments.
Figure 3: Efficiency of long-document models vs respective (truncation) FirstP baselines. The figure shows an average relative gain (in %) vs. relative increase in run-time compared to respective FirstP baselines on MS MARCO, TREC DL 2019-2021, and Robust04 (for a representative subset of models). Except LongP RankGPT, LongP models truncate documents to be at most 1431 tokens. There is no truncation for RankGPT.
B.2 Miscellaneous Ablations
B.2.1 Varying the Number of Chunks
@@
To understand if truncating input to at most 1431 BERT tokens biased our experimental outcomes, we carried out an ablation study where one of the top-performing models was trained and evaluated using inputs of varying maximum lengths. To this end, we used PARADE Attention with the number of input chunks varying from one to six. For each setting, the same number of chunks was used during training and evaluation, i.e., we had to carry out six experiments. Similar to our main experiments, we trained each model using three seeds. We carried out this ablation experiment using our MS MARCO and Robust04 datasets.
@@
The results are presented in Table 6. We can see that—depending on the dataset—three or four input chunks are optimal. However, compared to using three chunks, the additional gains over the FirstP baseline are at most 0.6% when averaged over all test sets.
@@
Gao and Callan 2022 carried out a similar ablation using ClueWeb09. Increasing the number of input chunks from three to six lead to only about 2.3% relative improvement in NDCG@20. However, even this modest gain could have been slightly inflated due to model not being trained directly on shorter inputs. Indeed, truncation of an input for a six-chunk model to one chunk is potentially less effective than training and evaluating the model using one-chunk data.
B.2.2 Efficiency Evaluation
We conducted an efficiency evaluation because long-document models, despite their potential to capture more context, often incur substantial computational overhead compared to their FirstP counterparts, raising questions about their practical utility.
[…]
All LongP models are even slower and show less improvement. Given such small benefits at the cost of a substantial slow-down, one could question practicality of such models and suggest using FirstP variants instead.
Appendix C Additional Dataset Details
C.1 Summary Dataset Statistics
# of queries
avg. # of
BERT tokens
avg. # of
pos. judgements
MS MARCO doc. v1
MS MARCO doc. train 352K 7 1
MS MARCO doc. dev 5193 7 1
TREC DL 2019 43 7 153.4
TREC DL 2020 45 7.4 39.3
MS MARCO v2
TREC DL 2021 57 9.8 143.9
Robust04
title 250 3.6 69.6
description 250 18.7 69.6
MS MARCO FarRelevant
train 50K 7.0 1
test 1K 7.0 1
LongEmbed
Needle 800 13.0 1
Passkey 800 9.7 1
MS MARCO pass. v1
TREC DL 2019 43 7 95.4
TREC DL 2020 54 7.2 66.8
BEIR
Natural Questions 3452 9.9 1.2
Touche 49 19 7.9
TREC COVID 50 493.5 16
NFCorpus (NFC) 323 38.2 5
DBP 400 38.2 6.5
SciFact 300 1.1 20.8
SciDocs 1000 4.9 13.1
Table 7: Query Statistics
data set # of documents average # of
BERT tokens
per document
Long-document collections
MS MARCO doc. v1 3.2M 1.4K
MS MARCO doc. v2 12M 2K
Robust04 0.5M 0.6K
MS MARCO FarRelevant 0.53M 1.1K
Needle (LongEmbed) 0.8K variable-length
Passkey (LongEmbed) 0.8K variable-length
Short-document collections
MS MARCO pass. v1 8.8M 75
Short-document collections: BEIR subset
Natural Questions 2.7M 107
Touche 382.5K 382
TREC COVID 171.3K 243
NFCorpus (NFC) 3.6K 355
DBP 4.6M 71
SciFact 5.2K 335
SciDocs 25.7K 234
  • LongEmbed subsets each have 16 subsets of documents whose lengths vary from (approximately) 256 to 32768 tokens.
Table 8: Document Statistics
Summary query and dataset statistics are shown in Tables 7 and 8. In the case of BEIR, we use seven subsets, which include:
  • Natural Questions Kwiatkowski et al. (2019);
  • Touché 2020 Bondarenko et al. (2020);
  • TREC COVID Voorhees et al. (2021);
  • NFCorpus Boteva et al. (2016);
  • DBPedia-Entity-v2 Hasibi et al. (2017);
  • SciFact Wadden et al. (2020);
  • SciDocs Cohan et al. (2020).
Please, note that in the case of MS MARCO FarRelevant, we created about 500K training and 7K testing queries, but to reduce experimentation cost we ended up using only 50K and 1K queries, respectively.
@@ C.1 Summary Dataset Statistics
FirstP (E5) 0.380 0.641 0.722 0.317
FirstP RankGPT (OpenAI) 0.708 0.790 0.352
FirstP RankGPT (Anthropic) 0.703 0.776 0.347
@@
LongP (E5) 0.353a $(-7.1\%)$ 0.649 $(+1.3\%)$ 0.724 $(+0.3\%)$ 0.323 $(+1.8\%)$
LongP RankGPT (OpenAI) 0.706 $(-0.3\%)$ 0.783 $(-1.0\%)$ 0.350 $(-0.7\%)$
LongP RankGPT (Anthropic) 0.707 $(+0.5\%)$ 0.780 $(+0.5\%)$ 0.348 $(+0.4\%)$
@@
Table 9: Comparison between long-document models and respective FirstP (truncation) baselines. Results for MS MARCO and TREC DL.
@@
FirstP RankGPT (OpenAI) 0.562 0.456 0.280
FirstP RankGPT (Anthropic) 0.541 0.446 0.268
@@
LongP RankGPT (OpenAI) 0.562 $(+0.0\%)$ 0.456 $(-0.0\%)$ 0.281 $(+0.3\%)$
LongP RankGPT (Anthropic) 0.538 $(-0.6\%)$ 0.445 $(-0.2\%)$ 0.268 $(-0.0\%)$